# TrialSynth: Generation of Synthetic Sequential Clinical Trial Data

### v1

@@ A.4 Baselines
PARSynthesizer from SDV (Zhang et al., 2022; Patki et al., 2016) since it is the most relevant model for synthesizing sequential event data, based on a conditional probabilistic auto-regressive (CPAR) model. To the best of our knowledge, no other models specifically handle sequential event data generation from scratch with easily accessible code.
TabDDPM (Kotelnikov et al., 2023) is a recently proposed state-of-the-art general tabular synthesizer based on diffusion models. Although it is not explicitly built for sequential data, we are able to enhance it by adding time as a numerical column. This model also outperforms CTGAN models Xu et al. (2019); Zhao et al. (2021, 2022), the previous go-to for synthetic tabular data generation. We believe that this is a strong, representative baseline of general tabular synthetic data generation.
HALO (Theodorou et al., 2023) is a hierarchical autoregressive language model that has achieved state-of-the-art performance for Electronic Health Record (EHR) synthesis. Still, it does not perform well on the clinical trial evaluation datasets, primarily due to the small size of training data, demonstrating the difficulty of this task.
A.5 TrialSynth Hyperparameters
For PARSyntheizer, default hyper-parameters were used. For TabDDPM, we followed the GitHub example for churn2 https://github.com/yandex-research/tab-ddpm, but trained for 10,000 steps for each dataset. Total running time took around 5-6 days on a NVIDIA 2080 GPU.
Table 6: Hyperparameters Considered for TrialSynth
Parameter Space
embedding_size [32,64,128]
patient_embedding_size [64,128,256]
num_transformer_layers (Encoder) [1,2,3,4,5,6,7,8]
num_heads (Encoder) [2,4,8]
num_transformer_layers (Decoder) [1,2,3,4,5,6,7,8]
num_heads (Decoder) [2,4,8]
lr [1e-3, 1e-4]
Table 7: Hyperparameters Considered for LSTM VAE
Parameter Space
embedding_size [32,64,128]
patient_embedding_size [64,128,256]
num_lstm_layers (Encoder) [1,2]
hidden_size (Encoder) [32,64,128]
num_lstm_layers (Decoder) [1,2]
hidden_size (Decoder) [32,64,128]
lr [1e-3, 1e-4]
@@ A.8 Ablations
Table 9: AUCROC results ($\uparrow$ higher the better, $\pm$ standard deviation) from fitting a downstream LSTM to predict death event from the ablation of informing of the exact length of the sequence to generate.
Dataset
Original
Data
LSTM
VAE
TrialSynth
(Events
Unknown)
TrialSynth
(Events
Known)
LC1 0.689±0.105 0.485±0.050 0.574±0.071 0.761±0.038
BC1 0.678±0.078 0.613±0.037 0.643±0.039 0.652±0.042
CC 0.637±0.140 0.403±0.094 0.632±0.017 0.654±0.098
BC2 0.670±0.128 0.560±0.053 0.557±0.057 0.618±0.053
BC3 0.632±0.072 0.437±0.040 0.589±0.041 0.624±0.037
VTE 0.640±0.038 0.496±0.021 0.585±0.022 0.642±0.022
LC2 0.738±0.149 0.608±0.092 0.717±0.083 0.860±0.056
Assuming Knowledge of Event Lengths In this section, we examine the ability of our framework to take in additional subject-level information regarding sequence generation. For example, TrialSynth assumes that we have access to the original event lengths (e.g., event 1 occurs 5 times in the original sequence, generate the times at which event 1 occurs), as well as the event indices (e.g., we know that subject 1 has events 2,5,6). For more details, see Section 3.3. Table 9 shows the result of allowing the model to know how many times events occur (for TrialSynth (Events Known)) or how the total length of the sequence to be generated (for TrialSynth (Events Unknown) and LSTM VAE). We see that providing the model with the event lengths boosts performance, as TrialSynth variants as well as the VAE LSTM in Table 9 performs better than in Table 2.
Table 10: Binary death event classification ROCAUCs ($\uparrow$ higher the better, $\pm$ standard deviation) of a downstream LSTM trained on data generated from the TrialSynth models comparing probabilistic decoding vs greedy decoding. Note that the sequence stopping length is the same for both approaches.
Dataset Original Data Probabilistic Greedy
LC1 0.689±0.105 0.572±0.051 0.453±0.052
BC1 0.678±0.078 0.646±0.037 0.629±0.042
CC 0.637±0.140 0.622±0.016 0.391±0.094
BC2 0.660±0.128 0.548±0.024 0.451±0.092
BC3 0.632±0.072 0.590±0.050 0.450±0.014
VTE 0.640±0.038 0.568±0.018 0.383±0.061
LC2 0.738±0.149 0.689±0.084 0.500±0.034
Autoregressive Greedy Sampling vs Probabilistic Sampling In this section, we compare the results of greedily choosing the best type as opposed to probabilistic sampling from the re-scaled logits $l\in\mathbb{R}^{N_{events}}$ when generating events from TrialSynth (Events Unknown). We re-scale the logits to be positive and sum to 1: $l=\frac{l-min(l)}{\sum l-min(l)}$, and sample from $k\sim Multinomial(l)$ (For the Greedy version, we take $\operatorname*{arg\,max}_{k}(l)$). Let Probabilistic represent the probabilistic sampling and Greedy represent greedy sampling. Table 10 shows the results, and we see that the Greedy version often fails to generate useful synthetic sequences.

### v2

@@ A.5 Baselines
PARSynthesizer from SDV [48, 33] since it is the most relevant model for synthesizing sequential event data, based on a conditional probabilistic auto-regressive (CPAR) model. To the best of our knowledge, no other models specifically handle sequential event data generation from scratch with easily accessible code.
TabDDPM [25] is a recently proposed state-of-the-art general tabular synthesizer based on diffusion models. Although it is not explicitly built for sequential data, we are able to enhance it by adding time as a numerical column. This model also outperforms CTGAN models [45, 49, 50], the previous go-to for synthetic tabular data generation. We believe that this is a strong, representative baseline of general tabular synthetic data generation.
HALO [36] is state-of-the art hierarchical autoregressive language model that has achieved state-of-the-art performance for Electronic Health Record (EHR) synthesis. Still, it does not perform well on the clinical trial evaluation datasets, primarily due to the small size of training data, demonstrating the difficulty of this task.
@@ A.8 Ablations
In this section, we include additional ablations on varying the multiplier on the standard deviation predicted by TrialSynth (Events Unknown), shown in Figure 7.
Figure 7: Additional Privacy / Utility tradeoffs examples in TrialSynth: Performance of distance to closest record (DCR) (red) and downstream ROC (blue) metrics at varying levels of VAE sampling variance (from 0.1 to 4), represented as the “Var Multiplier.”